LGSW FoV center on PatCams

10:00: spots are within the LGSW FoV, we have strong coma: we must flatten the shell. We close the loop on LGSW only and then we save a shell flat: 20150630_ARGOS_LGSW.sav.

HEXA -1.1 -7.9 -2.144 -51 245
BOARD -69.9 -47.6 39.3    
While the loop was closed on LGSW we take the best focus position on FLAO nulling the Truth sensing modes (FW1 on 50-50). Current positions are best for matching the 3 CalUnit spots within the 3 LGSW FoV. We must find the best position for each individual FoV:

Board X Board Y TN Fluxes
0 0 20150630_172142 5000, 4600, 4600
0.6 0 20150630_172601 5300, 4700, 200
0.9 0 - 5000, 0, 0
1.2 0 20150630_172740 4000, 0, 0
1.5 0 - 0, 0, 0
-0.6 0 20150630_172950 5800, 4000, 4500
-1.2 0 20150630_173109 0, 0, 3800
-1.8 0 20150630_173146 0, 0, 0
0 0.6 20150630_173327 3600, 1300, 4500
0 0.9 - 0, 0, 3800
0 1.2 - 0, 0, 1000
0 1.5 - 0, 0, 0
0 -0.3 - 5000, 4800, 2500
0 -0.6 - 5000, 4600, 0
0 -0.9 - 4200, 4600, 0
0 -1.2 - 1200, 4500, 0
0 -1.5 - 0, 3400, 0
0 -1.8 - 0, 0, 0
The offset of the best FoV center on the BlueChannel is [0.15; -0.3] wrt the common center position = [-69.75; 47.9] We set the stages there and then we move by +/-3.0mm from this position to get the spot position on the PatCam.

Board X Board Y TN PatCam pos
3.0 0.0 20150630_174759 [204; 127]
-3.0 0.0 20150630_175231 [156; 128]

Central pixel for the Blue PatCam is [180; 128]

We are also able to evaluate the patrol camera plate scale: 6mm = 10'' = 48px (@ bin4) = 0.208''/px

Comparing this value with the old one (taken by eye) we have an offset of [0.6; 0.4]''.

Yellow offset wrt P0 = [-0.12; -0.45] => Yellow center = [-70.0; -48.0]

Red offset wrt P0 = [-0.8; 0.9] => Red center = [-70.7; -46.7]

Board X Board Y TN PatCam pos
0.0 0.0 20150630_213449 spot is in LGSW
3.0 0.0 20150630_213706 [199; 107]
-3.0 0.0 20150630_213951 [158; 130]

Central pixel for the Yellow PatCam is [178; 118]

Board X Board Y TN PatCam pos
0.0 0.0 20150630_214304 spot is in LGSW
3.0 0.0 20150630_214531 [205; 128]
-3.0 0.0 20150630_214813 [158; 121]

Central pixel for the Red PatCam is [181; 124]

Red seems quite out wrt the old position... we try again the scan starting from [-70.7; -46.7] and we take the stage positions at which the Red pupil has 0 flux.

Board X Board Y Red flux Dir
-70.7 -46.7 2200 center
-70.7 -45.8 0 +Y
-70.7 -48.2 0 -Y
-70.7 -47.0 2200 0
-69.3 -47.0 0 +X
-71.8 -47.0 0 -X

DX= 2.5mm = 4.2"

DY= 2.4mm = 4.0"

Best FoV Center for RedChannel @ [-70.55; -47.0]

Board X Board Y TN PatCam pos
-73.55 -47.0 20150630_220453 [159; 123]
-67.55 -47.0 20150630_220759 [205; 130]

So new central pixel for the Red PatCam is [182; 126]

The 3 hot pixel positions have been saved into a new Patrol camera target: 20150701_075500.fits.

Piezo initial position

We go to the center of the Red FoV moving stages to [-70.7; -47.0]. Then we close the piezo loop and we take the mirror offset. We repeat the procedure for all the 3 channels:

Pupil PI X PI Y Pup X Pup Y
Blue -5.06 0.20 768 514
Yellow -2.00 3.20 270 722
Red -4.02 -0.61 473 663

These new piezo positions have been saved into a new jitter initial position: 2015701_080100.fits.

LUCI2 best focus (N30 and N3.75)

11:00: we check the best focus on LUCI using N30 camera and closing the loop on FLAO.

BIN1: REC= 20150125_161348, because of dome light we cannot use BIN1

BIN2: REC= 20130916_221257, we take images on N30 while sweeping the Z stage.

We check the old N30 images taken during the FLAO+LUCI commissioning that were getting almost 70% SR with BIN1: Z stage was at 55mm (bwt LUCI was removed and reinstalled in the meanwhile), rotator angle was 210deg. Reference file: luci2.20150404.0173.fits.

So we move the LUCI rotator to 210deg, we set the stage to 55mm and we optimize the FLAO gains: we have something really similar... We sweep between 50 and 60 and we find the best focus N30 = 56.5mm. We take LUCI images at each step in the end we take 20 dark frames.

We move to N3.75 and we find that the best focus N3.75 = 60mm

New Rec test

15:45: we try the new TM reconstructors:

REC Tag TN
TCoG48 bin2 20150630_225001
20150630_225032
GLAO1 pyr bin2 20150630_231404
20150630_231504

The 2 RECs seems to behave identically...

Mirror Panic @ 2015-06-30 23:16:34

Pockels cells test

17:00: at the end of the daytime we check the PC suppression:

TN Sup
'20150630_020735', '20150630_020902' 94, 113, 125

The Blue and Red cells shows worse performance than yesterday, double hear start to appear...

-- MarcoBonaglia - 01 Jul 2015
Topic revision: r3 - 01 Jul 2015, MarcoBonaglia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback